Theory and Applications of Categories, Vol. 5, No. 3, 1999, pp. 70–80.

A NOTE ON THE EXACT COMPLETION OF A REGULAR CATEGORY, AND ITS INFINITARY GENERALIZATIONS

STEPHEN LACK

Transmitted by Walter Tholen

ABSTRACT. A new description of the exact completion C_{ex/reg} of a regular category C is given, using a certain topos Shv(C) of sheaves on C; the exact completion is then constructed as the closure of C in Shv(C) under finite limits and coequalizers of equivalence relations. An infinitary generalization is proved, and the classical description of the exact completion is derived.

1. Introduction

A category C with finite limits is said to be regular if every morphism factorizes as a strong epimorphism followed by a monomorphism, and moreover the strong epimorphisms are stable under pullback; it follows that the strong epimorphisms are precisely the regular epimorphisms, namely those arrows which are the coequalizer of their kernel pair. Every kernel pair is an equivalence relation; a regular category is said to be exact if every equivalence relation is a kernel pair. Thus a regular category is a category with finite limits and coequalizers of kernel pairs, satisfying certain exactness conditions; while an exact category is a category with finite limits and coequalizers of equivalence relations, satisfying certain exactness conditions. Regular and exact categories were introduced by Barr [1], but the definition given here is due to Joyal.

A functor between regular categories is said to be regular if it preserves finite limits and strong (=regular) epimorphisms. There is a 2-category Reg of regular categories, regular functors, and natural transformations, and it has a full sub-2-category Ex comprising the exact categories. The inclusion of Ex into Reg has a left biadjoint, and the value of this biadjoint at a regular category C is what we mean by the exact completion of the regular category C.

Free regular and free exact categories have received a great deal of attention. A syntactic description of the free exact category on a category with finite limits was given by Carboni and Celia Magno in [2]. It was then observed that the same construction could be carried out starting with a category not with finite limits, but only with weak finite limits; this construction, along with its universal property, was described by Carboni and Vitale in [3]. The same paper also contained a description of the free regular category on a category with weak finite limits. Meanwhile, a quite different account of the free exact category on a finitely complete one was given by Hu in [10], namely as the full subcategory of the functor category [Lex(C, Set), Set] comprising those functors which preserve finite products and filtered colimits, where Lex(C, Set) is the category of finite-limit-preserving functors from C to Set. Finally, Hu and Tholen described in [11] the free exact category, and the free regular category, on a category C with weak finite limits, as full subcategories of the presheaf category [C^{op}, Set]: the free exact category was constructed as the closure of C in [C^{op}, Set] under coequalizers of equivalence relations, while the free regular category was constructed as the closure of C in [C^{op}, Set] under coequalizers of kernel pairs.

The support of the Australian Research Council is gratefully acknowledged.

Received by the editors 1998 November 6 and, in revised form, 1999 January 15.

Published on 1999 January 19.

1991 Mathematics Subject Classification: 18A35, 18A40, 18E10, 18F20.

Key words and phrases: regular category, exact category, exact completion, category of sheaves.

\begin{matrix}c \\ ◯ \end{matrix} Stephen Lack 1999. Permission to copy for private use granted.

In a different line, the exact completion of a regular category had already been described by Lawvere in [14], and more fully by R. Succi Cruciani [5]; see also the account of Freyd and Scedrov in [7]. It was also used in [3] in constructing the free exact category on a category with weak finite limits. The construction is conceptually attractive: one forms the (bi)category Rel(C) of relations in the regular category C, and then freely splits those idempotents which correspond to equivalence relations in C; this gives the (bi)category of relations, Rel(C_{ex/reg}), in the exact completion C_{ex/reg} of the original regular category, and one can extract C_{ex/reg} as the category of maps in the bicategory Rel(C_{ex/reg}), that is, as the category of those arrows which have a right adjoint in the bicategory.

The purpose of this note is to provide an alternative description of C_{ex/reg}, analogous to the description in [11] of the exact and regular completions of a category with weak finite limits, in which C_{ex/reg} is seen as a full subcategory of the presheaf category [C^{op}, Set].

A disadvantage of this approach is that it applies only to small regular categories, but in fact this is less serious than it might seem. When we speak of small sets, we mean sets whose cardinality is less than some inaccessible cardinal ∞. A small category is then a category with a small set of objects and small hom-sets. Given a regular category C which is not small, it will suffice to choose an inaccessible cardinal ∞′ greater than the cardinality of each of the hom-sets of C and that of the set of objects of C; for we may then write SET for the category of sets whose cardinality is less than ∞′, and redefine small to mean “ of cardinality less than ∞′”. The exact completion of C can now be constructed as a full subcategory of [C^{op}, SET], just as in Section 3 below.

A joint paper with Kelly, still in preparation, will describe the precise sense in which all of the constructions described above are free cocompletions with respect to certain classes of colimits “in the lex world”.

2. Sheaves on a regular category

Say that a diagram

\xymatrix{K \ar[r]^{k} \ar[r]_{l} & A \ar[r]^{p} & B}

in a regular category is an exact fork if p is the coequalizer of k and l, and (k, l) is the kernel pair of p.

We shall, as promised, construct the free exact category on the regular category C as a full subcategory of the presheaf category [C^{op}, Set], but it is not in fact the presheaf category with which we mostly work, for in passing from C to C_{ex/reg} we wish to preserve coequalizers of kernel pairs in C, and the Yoneda embedding does not preserve such coequalizers. We therefore consider only those presheaves F for which, if

\xymatrix{K \ar[r]^{k} \ar[r]_{l} & A \ar[r]^{p} & B}

is an exact fork in C, then any arrow u : yA → F which coequalizes k and l must factor uniquely through p. By the Yoneda lemma, these are precisely the presheaves F : C^{op} → Set which preserve the equalizers of those pairs which are in fact cokernel pairs. We write Shv(C) for the full subcategory of [C^{op}, Set] given by these presheaves; clearly it contains the representables, and moreover is reflective by [6, Theorem 5.2.1].

The reason for the name Shv(C) is of course that the objects of Shv(C) are the sheaves for a (Grothendieck) topology on C. Since C is regular, there is a pretopology [12] (or basis for a topology [15]) on C for which the covering families of B are the singleton families (p : A → B) with p a strong epimorphism. This generates a topology on C, called the regular epimorphism topology [1, Section I.4], for which the sheaves are precisely the objects of Shv(C). We conclude that Shv(C) is a Grothendieck topos, and in particular an exact category. The inclusion Y : C → Shv(C) preserves whatever limits exist in C, since the Yoneda embedding preserves any limits which exist in C, and the inclusion of Shv(C) in the presheaf category [C^{op}, Set] preserves and reflects all limits. Thus Y : C → Shv(C) preserves finite limits; by construction it preserves coequalizers of kernel pairs, and so is a regular functor. This full regular embedding of C in the topos Shv(C), was described already in [1], as was the reflectivity of Shv(C) in [C^{op}, Set].

We shall need the following characterization of strong epimorphisms in Shv(C) as “local surjections”; since it is a (well-known) special case of [15, Corollary III.7.5], we only sketch the proof.

2.1. Lemma. An arrow α : F → G in Shv(C) is a strong epimorphism if and only if for every v : Y B → G, there exists a strong epimorphism \xymatrix{p : A \ar@{->>}[r] & B} in C and a factorization v.Y p = α.u.

Proof. Suppose that α is strong epi; for each object B of C, define HB to be the set \{x ∈ GB | (Gp)x = (αA)w for some w ∈ F A and some strong epi \xymatrix{p : A \ar@{->>}[r] & B} in C\}, and write mB : HB → GB for the inclusion. This is easily seen to be functorial in B, giving a functor H : C^{op} → Set, and a natural transformation m : H → G all of whose components are monomorphisms; clearly α factorizes through m. One now shows that H is in fact a sheaf, so that the strong epimorphism α in Shv(C) factors through the subobject m, which must therefore be invertible. It follows that α satisfies the conditions of the lemma.

Suppose, conversely, that α : F → G satisfies the conditions of the lemma. If α factorizes (in Shv(C)) as α = m.β with m : H → G monic, then for any v : Y B → G we have a commutative diagram

[figure: Y A; Y p; u; Y B; v; F; β; H; m; G]

in Shv(C) with Y p strong epi and m mono, and so a factorization of v through m. Since the representables form a strong generator for Shv(C), it follows that m is invertible, and so that α is a strong epimorphism.

We also need the following description of geometric morphisms into Shv(C); once again it is a special case of a well-known result.

2.2. Lemma. Let E be a Grothendieck topos. Composition with Y : C → Shv(C) induces an equivalence between Reg (C, E) and the category Geom (E, Shv(C)) of geometric morphisms from E to Shv(C); the inverse image S^{∗} of the geometric morphism corresponding to S : C → E is given by the (pointwise) left Kan extension Lan_{Y} S of S along Y : C → Shv(C).

Proof. There is a well-known equivalence between Geom(E,[C^{op}, Set]) and the category Lex(C,E) of finite-limit-preserving functors from C to E; it is a special case, for example, of [15, Theorem VII.7.2]. By a special case of [15, Lemma VII.7.3], a geometric morphism from E to [C^{op}, Set] factors through Shv(C) if and only if the corresponding finite-limit-preserving functor S : C → E maps strong epimorphisms in C to epimorphisms in E; but, since E is a topos, all epimorphisms are strong, and so this just says that S is a regular functor.

Finally the inverse image S^{∗} of the geometric morphism corresponding to a regular functor S : C → E satisfies S^{∗}L ≅ Lan_{JY} S, where L : [C^{op}, Set] → Shv(C) is the reflection, and so S^{∗} ≅ S^{∗}LJ ≅ (Lan_{JY} S)J ≅ (Lan_{J}Lan_{Y} S)J ≅ Lan_{Y} S.

3. The exact completion of a regular category

We now write C_{ex/reg} for the full subcategory of Shv(C) comprising those objects which are coequalizers in Shv(C) of equivalence relations in C. We write Z : C ↪ C_{ex/reg} and W : C_{ex/reg} ↪ Shv(C) for the inclusions.

3.1. Lemma. If \xymatrix{p : Y A \ar@{->>}[r] & F} is a strong epimorphism and m_{i} : F → G_{i} is a finite jointly-monic family in Shv(C), then F lies in C if each G_{i} does; thus in particular C is closed in C_{ex/reg} under subobjects.

Proof. Since p is strong epi, it is the coequalizer of its kernel pair. Since the m_{i} are jointly monic, the induced arrow m : F → Π_{i}G_{i} is monic, and so the kernel pair of p is also the kernel pair of mp. If each G_{i} is in C, then since C has finite products, Π_{i}G_{i} is in C; and since C has pullbacks, the kernel pair of mf lies in C; and so F lies in C since C has coequalizers of kernel pairs. The second statement follows from the first: if F lies in C_{ex/reg} then there is a strong epimorphism \xymatrix{Y A \ar@{->>}[r] & F} ; if also F is a subobject of some Y B then we conclude that F lies in C.

3.2. Proposition. C_{ex/reg} is closed in Shv(C) under finite limits and coequalizers of equivalence relations; thus C_{ex/reg} is an exact category, the inclusions Z : C ↪ C_{ex/reg} and W : C_{ex/reg} ↪ Shv(C) are fully faithful regular functors, and C_{ex/reg} is the closure of C in Shv(C) under finite limits and coequalizers of equivalence relations.

Proof. Shv(C) is exact, the inclusion of C into Shv(C) is regular, and every object of C_{ex/reg} is the coequalizer in Shv(C) of an equivalence relation in C; thus it will suffice to prove the first statement: that C_{ex/reg} is closed under finite limits and coequalizers of equivalence relations.

Step 1: C_{ex/reg} is closed in Shv(C) under finite products. Of course the terminal object of Shv(C) lies not just in C_{ex/reg} but in C. Given a finite non-empty family (F_{i})_{i∈I} of objects of C_{ex/reg}, we have exact forks

\xymatrix{Y R_{i} \ar[r] \ar[r] & Y A_{i} \ar[r]^{p} & F_{i}}

and so, since Shv(C) is exact, an exact fork

\xymatrix{Π_{i}Y R_{i} \ar[r] \ar[r] & Π_{i}Y A_{i} \ar[r] & Π_{i}F_{i} ;}

but C has finite products, and so Π_{i}Y R_{i} and Π_{i}Y A_{i} are both in C, whence Π_{i}F_{i} is in C_{ex/reg}.

Step 2: If

\xymatrix{P \ar[r]^{p_{2}} \ar[d]_{p_{1}} & Y B \ar[d]^{v} \\ Y A \ar[r]_{u} & G}

is a pullback in Shv(C), and G is in C_{ex/reg}, then P is in C. For since G is in C_{ex/reg}, we have an exact fork

\xymatrix{Y R \ar[r]^{r} \ar[r]_{s} & Y C \ar[r]^{p} & G}

in Shv(C), and so, by Lemma 2.1, a diagram

\xymatrix{Y A′ \ar[r]^{u′} \ar@{->>}[d]_{a} & Y C \ar@{->>}[d]^{p} & Y B′ \ar[l]_{v′} \ar@{->>}[d]^{b} \\ Y A \ar[r]_{u} & G & Y B \ar[l]^{v}}

with a and b strong epis.

We can therefore form diagrams of pullbacks

\xymatrix{Y D \ar[r] \ar[d] & Y R_{B} \ar[r] \ar[d] & Y B′ \ar[d]^{v′} & Y D \ar@{->>}[r]^{a′′} \ar@{->>}[d]_{b′′} & P_{B} \ar[r] \ar@{->>}[d]^{b′} & Y B′ \ar@{->>}[d]^{b} \\ Y R_{A} \ar[r] \ar[d] & Y R \ar[r]^{s} \ar[d]_{r} & Y C \ar[d]^{p} & P_{A} \ar@{->>}[r]_{a′} \ar[d] & P \ar[r]^{p_{2}} \ar[d]_{p_{1}} & Y B \ar[d]^{v} \\ Y A′ \ar[r]_{u′} & Y C \ar[r]_{p} & G & Y A′ \ar@{->>}[r]_{a} & Y A \ar[r]_{u} & G}

in Shv(C), giving

[figure: Y D; b′a′′; P; ^{\binom{p_{1}}{p_{2}}}; Y A × Y B ;]

and now P lies in C by Lemma 3.1.

Step 3: C_{ex/reg} is closed in Shv(C) under pullbacks. If

\xymatrix{P \ar[r]^{p_{2}} \ar[d]_{p_{1}} & F′ \ar[d]^{v} \\ F \ar[r]_{u} & G}

is a pullback in Shv(C) with F, F′, and G in C_{ex/reg}, we have strong epis \xymatrix{q : Y A \ar@{->>}[r] & F} and \xymatrix{q′ : Y A′ \ar@{->>}[r] & F′} , and so a diagram

\xymatrix{Q \ar@{->>}[r]^{q_{2}} \ar@{->>}[d]_{q′_{2}} & P′_{1} \ar[r] \ar@{->>}[d]^{q′_{1}} & Y A′ \ar@{->>}[d]^{q′} \\ P_{1} \ar@{->>}[r]_{q_{1}} \ar[d] & P \ar[d]_{p_{1}} \ar[r]^{p_{2}} & F′ \ar[d]^{v} \\ Y A \ar@{->>}[r]_{q} & F \ar[r]_{u} & G}

of pullbacks in Shv(C). Thus we have

[figure: Q; q′_{1}q_{2}; P; ^{\binom{p_{1}}{p_{2}}}; F × F′]

with Q in C by Step 2, and F × F′ in C_{ex/reg} by Step 1. Now q′_{1}q_{2} is strong epi, and so must be the coequalizer of its kernel pair. But \binom{p_{1}}{p_{2}} is monic, and so the kernel pair of q′_{1}q_{2} is the kernel pair of \binom{p_{1}}{p_{2}}q′_{1}q_{2}, which, by Step 2, must lie in C. We conclude that P is the coequalizer of an equivalence relation in C, and so lies in C_{ex/reg}.

Step 4: C_{ex/reg} is closed in Shv(C) under coequalizers of equivalence relations. Let

\xymatrix{F \ar[r]^{r} \ar[r]_{s} & G \ar[r]^{q} & H}

be an exact fork in Shv(C), with both F and G in C_{ex/reg}. Then we have a diagram

\xymatrix{P \ar@{->>}[r]^{p′_{1}} \ar@{->>}[d]_{p′_{2}} & F_{B} \ar@{->>}[r]^{s′} \ar@{->>}[d]_{p_{2}} & Y B \ar@{->>}[d]^{p} \\ F_{A} \ar@{->>}[r]^{p_{1}} \ar@{->>}[d]_{r′} & F \ar@{->>}[r]^{s} \ar@{->>}[d]_{r} & G \ar@{->>}[d]^{q} \\ Y B \ar@{->>}[r]_{p} & G \ar@{->>}[r]_{q} & H}

of pullbacks in Shv(C), in which Y B, F, and G are in C_{ex/reg}, and so F_{A} and F_{B} are in C_{ex/reg}, and so P is in C_{ex/reg}. Thus there is a strong epi \xymatrix{q′ : Y D \ar@{->>}[r] & P} ; but also there is a mono \binom{s′p′_{1}}{r′p′_{2}} : P → Y B × Y B, and so, by Lemma 3.1, we conclude that P is in C. Finally

\xymatrix{P \ar[r]^{s′p′_{1}} \ar[r]_{r′p′_{2}} & Y B \ar[r]^{qp} & H}

is an exact fork in Shv(C) with P and Y B in C, and so we conclude that H is in C_{ex/reg}.

We now prove the universal property of C_{ex/reg}; the proof is reminiscent of [13]:

3.3. Theorem. If D is an exact category, then composition with Z : C → C_{ex/reg} induces an equivalence of categories Reg(Z, D) : Reg(C_{ex/reg}, D) → Reg(C, D) to which the inverse equivalence is given by (pointwise) left Kan extension along Z; thus C_{ex/reg} is the exact completion of C.

Proof. Let S : C → D be a regular functor; then the composite Y S : C → Shv(D) of S with Y : D → Shv(D) is regular, and so by Lemma 2.2 induces a geometric morphism from Shv(D) to Shv(C), whose inverse image we call S^{∗} : Shv(C) → Shv(D).

Consider now the full subcategory C_{1} of Shv(C) containing those objects F for which S^{∗}F lies in D; it is the pseudopullback of Y : D → Shv(D) along S^{∗}. Since S^{∗} preserves finite limits and coequalizers of equivalence relations, and since D is closed in Shv(D) under finite limits and coequalizers of equivalence relations, it follows that C_{1} is closed in Shv(C) under finite limits and coequalizers of equivalence relations. Clearly C_{1} contains C, and so, by Proposition 3.2, C_{1} must contain C_{ex/reg}; thus S^{∗} restricts to a functor S̄ : C_{ex/reg} → D. Since S^{∗} and W preserve finite limits and coequalizers of equivalence relations, and since Y : D → Shv(D) reflects them, we deduce that S̄ is a regular functor. (It follows that Reg(Z, D) : Reg(C_{ex/reg}, D) → Reg(C, D) is essentially surjective on objects.)

By Lemma 2.2 we have S^{∗} ≅ Lan_{Y} (Y S), and so Y S̄ ≅ S^{∗}W ≅ Lan_{Y} (Y S)W ≅ (Lan_{W} Lan_{Z}(Y S))W ≅ Lan_{Z}(Y S). This says precisely that C_{ex/reg}(Z, F) ∗ Y S ≅ Y S̄F for each object F of C_{ex/reg}, where C_{ex/reg}(Z, F) ∗ Y S denotes the colimit of Y S : C → Shv(D), weighted by the functor C_{ex/reg}(Z, F) : C^{op} → Set taking an object C of C to C_{ex/reg}(ZC, F). But Y is fully faithful and so reflects colimits, giving C(Z, F) ∗ S ≅ S̄F; which is to say that Lan_{Z}(S) ≅ S̄.

Thus we obtain a functor Lan_{Z} : Reg(C, D) → Reg(C_{ex/reg}, D), with the composite Reg(Z, D).Lan_{Z} naturally isomorphic to the identity, since Z is fully faithful. It remains to show that the canonical map η : Lan_{Z}(T Z) → T is invertible, for any regular functor T : C_{ex/reg} → D. Now Lan_{Z}(T Z) and T both preserve coequalizers of equivalence relations, and ηC is invertible for objects of C, hence by Proposition 3.2 must be invertible for all objects of C_{ex/reg}.

4. Infinitary generalizations

In [11], the more general case of κ-regular and κ-exact categories was considered, for a regular cardinal κ. We recall that a κ-regular category is a regular category with κproducts, such that the strong epimorphisms are closed under κ-products; and that a κ-exact category is a κ-regular category that is exact. These classes of categories were introduced by Makkai [16] and further studied by Hu [9]; in the additive context, κ-regular categories were already considered by Grothendieck in [8].

We write κ-Reg for the 2-category of κ-regular categories, functors preserving κ-limits and strong epimorphisms, and natural transformations; and we write κ-Ex for the full sub-2-category of κ-Reg comprising the κ-exact categories. The value at a κ-regular category C of a left biadjoint to the inclusion of κ-Ex in κ-Reg is called the κ-exact completion of the κ-regular category C.

If C is a κ-regular category, and (F_{i})_{i∈I} is a κ-small family of objects of C_{ex/reg}, then for each i there is an exact fork

\xymatrix{Y R_{i} \ar[r] \ar[r] & Y A_{i} \ar[r] & F_{i}}

in Shv(C), and so, since Shv(C) is κ-exact, an exact fork

\xymatrix{Π_{i}Y R_{i} \ar[r] \ar[r] & Π_{i}Y A_{i} \ar[r] & Π_{i}F_{i}}

in Shv(C); but since C has κ-products, it follows that Π_{i}Y R_{i} and Π_{i}Y A_{i} are in C, and so that Π_{i}F_{i} is in C_{ex/reg}. Thus C_{ex/reg} is closed in Shv(C) under κ-products, and so is κ-exact, since Shv(C) is so. Thus C_{ex/reg} is a κ-exact category, and Z : C → C_{ex/reg} is a κ-regular functor.

Moreover, if D is a κ-exact category, and S : C → D a κ-exact functor, then it is clear by the construction of κ-products in C_{ex/reg} that Lan_{Z}S : C_{ex/reg} → D preserves κ-products, and so we have:

4.1. Theorem. If C is κ-regular, and D is κ-exact, then composition with Z : C → C_{ex/reg} induces an equivalence of categories κ-Reg(Z, D) : κ-Reg(C_{ex/reg}, D) → κ-Reg(C, D) to which the inverse equivalence is given by (pointwise) left Kan extension along Z; thus C_{ex/reg} is the κ-exact completion of C.

5. A concrete description of the exact completion

We can now apply Theorem 3.3 to derive a concrete description of the exact completion, using the language of relations in a regular category. For further details on the calculus of relations see [7, 4]; in fact all that we need is summarized in [3]. We write R : A ↛ B for a relation from A to B, by which we mean an equivalence class of subobjects of A × B. We further write S ∘R : A ↛ C for the composite of relations R : A ↛ B and S : B ↛ C, and R^{∘} : B ↛ A for the opposite relation of R. Recall that an equivalence relation is a relation R satsifying R ∘ R = R, R = R^{∘}, and 1 ≤ R.

We shall construct a category E equipped with a functor Φ : E → C_{ex/reg} which is fully faithful and essentially surjective on objects, and so an equivalence. We take the objects of E to be the equivalence relations in C, and define Φ on objects to take such an equivalence relation to its coequalizer in C_{ex/reg}. Since every object of C_{ex/reg} is the coequalizer of an equivalence relation in C, the functor Φ, once defined, will be essentially surjective on objects.

Given an equivalence relation R : A ↛ A, recall that the coequalizer r : A → A/R of R satisfies r^{∘} ∘ r = R and r ∘ r^{∘} = 1, and so r ⊣ r^{∘}. If S : B ↛ B is another equivalence relation, and s : B → B/S is its coequalizer, an arrow in C_{ex/reg} from A/R to B/S determines a relation from A to B as follows. We form the pullback

\xymatrix{D \ar[rr]^{f} \ar@{->>}[d]_{g} &  & B \ar@{->>}[d]^{s} \\ A \ar@{->>}[r]_{r} & A/R \ar[r]_{α} & B/S}

in C_{ex/reg}; of course this is a relation in C_{ex/reg}, but since D is a subobject of A × B, it follows by Lemma 3.1 that D is in fact in C, and so f and g are the legs of a relation from A to B in C. Since g and r are strong epimorphisms, there is clearly at most one α : A/R → B/S giving rise in this way to a particular relation U : A ↛ B, but not every U does so arise.

To say that a relation U : A ↛ B arises, as in the previous paragraph, from α : A/R → B/S, is precisely to say that U = s^{∘} ∘ α ∘ r. Since s ∘ s^{∘} = 1 and r ∘ r^{∘} = 1, it then follows that α = s ∘ s^{∘} ∘ α ∘ r ∘ r^{∘} = s ∘ U ∘ r^{∘}. Thus U arises from a (necessarily unique) α if and only if U = s^{∘} ∘ s ∘ U ∘ r^{∘} ∘ r and s ∘ U ∘ r^{∘} is a map (rather than a general relation).

The first condition says that U = S ∘ U ∘ R, which, since S and R are idempotents, is equivalent to the two conditions:

U = S ∘ U
U = U ∘ R.

It remains to express the condition that α = s ∘ U ∘ r^{∘} be a map; that is, that α ∘ α^{∘} ≤ 1 and 1 ≤ α^{∘} ∘ α. Now α ∘ α^{∘} = (s ∘ U ∘ r^{∘}) ∘ (s ∘ U ∘ r^{∘})^{∘} = s ∘ U ∘ r^{∘} ∘ r ∘ U^{∘} ∘ s^{∘} = s ∘ U ∘ R ∘ U^{∘} ∘ s^{∘} = s ∘ U ∘ U^{∘} ∘ s^{∘}; and, since s ⊣ s^{∘}, the inequality s ∘ U ∘ U^{∘} ∘ s^{∘} ≤ 1 is equivalent to U ∘ U^{∘} ∘ s^{∘} ≤ s^{∘} and so to U ∘ U^{∘} ≤ s^{∘} ∘ s = S; whence α ∘ α^{∘} ≤ 1 is equivalent to

U ∘ U^{∘} ≤ S.

On the other hand α^{∘} ∘ α = (s ∘ U ∘ r^{∘})^{∘} ∘ (s ∘ U ∘ r^{∘}) = r ∘ U^{∘} ∘ s^{∘} ∘ s ∘ U ∘ r^{∘} = r ∘ U^{∘} ∘ S ∘ U ∘ r^{∘} = r ∘ U^{∘} ∘ U ∘ r^{∘}; and 1 ≤ r ∘ U^{∘} ∘ U ∘ r^{∘} implies that R = r^{∘} ∘ r ≤ r^{∘} ∘ r ∘ U^{∘} ∘ U ∘ r^{∘} ∘ r = R^{∘} ∘ U^{∘} ∘ U ∘ R = (U ∘ R)^{∘} ∘ (U ∘ R) = U^{∘} ∘ U, while R ≤ U^{∘} ∘ U implies that 1 = r ∘ r^{∘} ∘ r ∘ r^{∘} = r ∘ R ∘ r^{∘} ≤ r ∘ U^{∘} ∘ U ∘ r^{∘}; and so 1 ≤ α^{∘}α is equivalent to

R ≤ U^{∘} ∘ U.

We now define an arrow in E from R to S to be a relation U : A ↛ B in C satisfying the three conditions S ∘ U ∘ R = U, U ∘ U^{∘} ≤ S, and R ≤ U^{∘} ∘ U; and define Φ : E(R, S) → C_{ex/reg}(A/R, B/S) be the bijection taking U to s ∘ U ∘ r^{∘}.

Composition in C_{ex/reg} now induces a composition in E making E into a category and Φ an equivalence of categories from E to C_{ex/reg}; to complete the description of E it remains only to calculate the induced composition and identities.

Suppose then that T : C ↛ C is another equivalence relation in C, and t : C → C/T is its coequalizer. Let β : B/S → C/T be an arrow in C_{ex/reg}, and V = t^{∘} ∘ β ∘ s the corresponding relation. Then V ∘U = t^{∘}∘β∘s∘s^{∘}∘α∘r = t^{∘}∘β∘α∘r, which is the relation corresponding to βα; thus composition in E is just composition of relations. On the other hand, the identity arrow 1 : A/R → A/R corresponds to the relation R : A ↛ A, which is therefore the identity arrow in E at the object R. Thus E is now seen to be precisely the exact completion as described in [3].

References

[1] Michael Barr, Exact categories, in Exact Categories and Categories of Sheaves, Lecture Notes in Mathematics 236, Springer-Verlag, 1971.

[2] A. Carboni and R. Celia Magno, The free exact category on a left exact one, J. Austral. Math. Soc. (Ser. A), 33:295–301, 1982.

[3] A. Carboni and E.M. Vitale, Regular and exact completions, J. Pure Appl. Alg., 125:79–116, 1998.

[4] A. Carboni and R.F.C. Walters, Cartesian Bicategories I, J. Pure Appl. Alg., 49:11–32, 1987.

[5] R. Succi Cruciani, La teoria delle relazioni nello studio di categorie regolari e categorie esatte, Riv. Mat. Univ. Parma, 4:143–158, 1975.

[6] P. Freyd and G.M. Kelly, Categories of continuous functors I, J. Pure Appl. Alg., 2:169–191, 1972.

[7] P. Freyd and A. Scedrov, Categories, Allegories, North-Holland Press, Amsterdam, 1990.

[8] A. Grothendieck, Sur quelques points d’algébre homologique, Tôhoku Math. J., 9:119–221, 1957.

[9] Hongde Hu, Dualities for accessible categories, Can. Math. Soc. Conf. Proc. 13:211–242, 1992.

[10] Hongde Hu, Flat functors and free exact categories, J. Austral. Math. Soc. (Ser. A), 60:143–156, 1996.

[11] Hongde Hu and Walter Tholen, A note on free regular and exact completions and their infinitary generalizations, Theory and Applications of Categories, 2:113–132, 1996.

[12] P.T. Johnstone, Topos Theory, L.M.S Monographs No. 10, Academic Press, London, 1977.

[13] G.M. Kelly, On the essentially-algebraic theory generated by a sketch, Bull. Austral. Math. Soc. 26:45–56, 1982.

[14] F.W. Lawvere, Category theory over a base topos (the “Perugia notes”), unpublished manuscript, 1973.

[15] Saunders Mac Lane and Ieke Moerdijk, Sheaves in Geometry and Logic: A First Introduction to Topos Theory, Springer-Verlag, 1992.

[16] M. Makkai, A theorem on Barr-exact categories, with an infinitary generalization, Ann. Pure Appl. Logic 47:225–268, 1990.

School of Mathematics and Statistics University of Sydney Sydney NSW 2006 AUSTRALIA

Email: stevel@maths.usyd.edu.au

This article may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anony-

mous ftp at ftp://ftp.tac.mta.ca/pub/tac/html/volumes/1999/n3/n3.{dvi,ps}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi and Postscript format. Details will be e-mailed to new subscribers and are available by WWW/ftp. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at URL http://www.tac.mta.ca/tac/ or by anonymous ftp from ftp.tac.mta.ca in the directory pub/tac/info. You may also write to tac@mta.ca to receive details by e-mail.

Editorial board.

John Baez, University of California, Riverside: baez@math.ucr.edu Michael Barr, McGill University: barr@math.mcgill.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Jean-Luc Brylinski, Pennsylvania State University: jlb@math.psu.edu Aurelio Carboni, Università dell Insubria: carboni@fis.unico.it P. T. Johnstone, University of Cambridge: ptj@pmms.cam.ac.uk G. Max Kelly, University of Sydney: kelly m@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.ruu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mscs.dal.ca Andrew Pitts, University of Cambridge: ap@cl.cam.ac.uk Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz James Stasheff, University of North Carolina: jds@charlie.math.unc.edu Ross Street, Macquarie University: street@macadam.mpce.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Sydney: walters b@maths.usyd.edu.au R. J. Wood, Dalhousie University: rjwood@mscs.dal.ca